An electron near the surface of the Earth has a velocity of 2,500 m/s in the +y -direction, through a region where there is a 2T magnetic field in the +x -direction & a 300 V/m electric field in the –z-direction (with upward defined as the +z -direction ).

a) What is the electric force on the electron?

b) What is the magnetic force on the electron?
Answer #1

given

speed v = 2500 m/s j

magnetic field B = 2T i

electric field E = 300 V/m (-k)

a) electric force Fe = qE = -1.6*10^-19*(-300) = 4.8*10^-17 k N (positive z direction)

b) magnetic force FB = q(v*B) = -1.6*10^-19(2500i*2j) = -8*10^-16 k N (negative z direction)

FB = 1.